OpenStudy (anonymous):

please help me >.< 1) 5 (11root) - 12 (11root) - 2(11root) = 2) (8 [6root] - [2root]) + (5 [6root] - 6 [2root]) =

jimthompson5910 (jim_thompson5910):

# 1 is \[\large 5\sqrt{11} - 12\sqrt{11} - 2\sqrt{11}\] right?

OpenStudy (anonymous):

yes exactly ! (: the answer i got was -9\[\sqrt{33}\]

jimthompson5910 (jim_thompson5910):

the -9 is correct, but the 33 is not

jimthompson5910 (jim_thompson5910):

think of x = sqrt(11) to get 5x - 12x - 2x combining all that gives you 5x - 12x - 2x = -9x now replace x with sqrt(11) to get -9*sqrt(11)

OpenStudy (anonymous):

-9 \[\sqrt{11}\]

OpenStudy (anonymous):

oooh thank you !

jimthompson5910 (jim_thompson5910):

yes exactly, yw

OpenStudy (anonymous):

but for the second question that method wouldnt work??

jimthompson5910 (jim_thompson5910):

think of x = sqrt(6) and y = sqrt(2)

jimthompson5910 (jim_thompson5910):

does that help?

OpenStudy (anonymous):

ummm.. alitle bit >.<

jimthompson5910 (jim_thompson5910):

(8*sqrt(6) - sqrt(2)) + (5*sqrt(6) - 6*sqrt(2)) (8x - y) + (5x - 6y) how about now?

OpenStudy (anonymous):

so then i would be (13x - 5y) ??

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

-y - 6y isn't -5y though

OpenStudy (anonymous):

oooh -7y

jimthompson5910 (jim_thompson5910):

so 13x - 7y

jimthompson5910 (jim_thompson5910):

now replace x with sqrt(6) and replace y with sqrt(2)

OpenStudy (anonymous):

i get it (: thank you soo much !!

jimthompson5910 (jim_thompson5910):

you're welcome